OpenStudy (anonymous):

The expression (1/50)(sqrt(1/50)+sqrt(2/50)+sqrt(3/50)+...+sqrt(50/50) is a Riemann Sum approximation for They want you to find its integral equivalent

OpenStudy (anonymous):

Well set up its integral equivalent, anyways

OpenStudy (anonymous):

It was multiple choice, but my teacher says that we need to know how to do all of them as if we weren't given the answer choices

OpenStudy (freckles):

is there no n's in your sum thingy above?

OpenStudy (anonymous):

Nope. They want you to set up a definite integral with it

OpenStudy (anonymous):

I know what the answer is, but I have no idea how to get to it

ganeshie8 (ganeshie8):

does this work ?\[\int\limits_0^1 \sqrt{x}~dx\]

OpenStudy (freckles):

oh approximation

OpenStudy (anonymous):

Yes, thats exactly it

ganeshie8 (ganeshie8):

the given riemann sum looks like "Upper sum with 50 intervals"

OpenStudy (anonymous):

How did you do that

OpenStudy (anonymous):

Whoooa

OpenStudy (freckles):

\[\frac{b-a}{n}=\Delta x \\ \text{ choosing } b=1 \text{ and } a=0 \text{ we have } \frac{1}{n} \\ \frac{1}{n} \cdot \sum_{i=1}^n \sqrt{0+i \cdot \frac{1}{n}} \\ \lim_{n \rightarrow \infty} \sum_{i=1}^n \sqrt{\frac{i}{n}} \frac{1}{n} =? \]

OpenStudy (anonymous):

Is n the number of subintervals?

OpenStudy (freckles):

yeah

OpenStudy (anonymous):

What is i?

OpenStudy (freckles):

it is good to memorize that \[\lim_{n \rightarrow \infty}\sum_{i=1}^{n} f(x_i) \Delta x=\int\limits_{a}^{b}f(x) dx \\ \text{ where } x_i=a+i \cdot \Delta x \\ \text{ and } \Delta x =\frac{b-a}{n}\]

OpenStudy (anonymous):

There is no way that I will be able to memorize that

OpenStudy (freckles):

\[\sum_{i=1}^{n}f(x_i) \Delta x\\ \Delta x(f(x_1)+f(x_2)+f(x_3) \cdots +f(x_n))\]

OpenStudy (anonymous):

oh my god

OpenStudy (freckles):

it is just a whole bunch of areas for rectangles being adding

OpenStudy (freckles):

i guess i can see how it looks way more scarier than that

OpenStudy (anonymous):

Why is it just sqrt x inside of the integral?

OpenStudy (freckles):

replace f(xi) with f(x) your xi is a+i delta x your xi is 0+i*1/n in this case

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wait...no...

OpenStudy (anonymous):

actualy yayayaya sorry :D

OpenStudy (freckles):

\[\int\limits_{a}^{b}f(x) dx=\lim_{n \rightarrow \infty} \sum_{i=1}^n f(x_i) \Delta x \\ \text{ where} \Delta x=\frac{b-a}{n} \text{ and } x_i=a+i \Delta x \\ \int\limits_a^b f(x) dx= \lim_{n \rightarrow \infty}\sum_{i=1}^{n}f(a+i \Delta x) \Delta x \\ \text{ Example } \lim_{n \rightarrow \infty} \sum_{i=1}^{n} f(5+i \frac{10}{n}) \cdot \frac{10}{n} \\ \frac{b-a}{n}=\frac{10}{n} \\ \text{ and we know } a=5 \text{ so we can find } b \\ \frac{b-5}{n}=\frac{10}{n} \\b=15 \\ \text{ so we have } \int\limits_5^{15} f(x) dx=\lim_{n \rightarrow \infty}\sum_{i=1}^{n}f(5+i \frac{10}{n}) \cdot \frac{10}{n}\]

OpenStudy (freckles):

so if f was sin(x) you would put sin(x)

OpenStudy (freckles):

or if f was kitty(x) you would put kitty(x)

ganeshie8 (ganeshie8):

http://gyazo.com/57bc0b924b5b9df6d6c99c55a597ea94

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty}\sum_{i=1}^{n}kitty(5+i \cdot \frac{10}{n}) \cdot \frac{10}{n}=\int\limits_{5}^{15} kitty(x) dx\]

OpenStudy (freckles):

where kitty is a mind boggling function of x

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

here is a more nice animation

OpenStudy (dan815):

lol did u make that

ganeshie8 (ganeshie8):

yes haha just testing what all geogebra can do :P

OpenStudy (dan815):

haha looks like u are having fun